# Computing the Intrinsic Delaunay Triangulation of a Closed Polyhedral Surface

### v1

@@ C.1 Analysis of InsertVertices
Proof C.3.
Let $m_{1}$ and $m_{2}$ count respectively the edges and the faces of $T^{1}$, and let $b$ count the boundary components of $\mathcal{S}(T)$. Double counting gives $3m_{2}\leq 2m_{1}$. Euler’s formula gives $m_{1}-m_{2}=m+2g+b-2$. And we have $b\leq m$. Therefore $m_{1}\leq 3m_{1}-3m_{2}<6(m+g)$.
Proof C.4 (Proof of Lemma C.1).
There are no more vertices inserted than there are edges in $T^{1}$, and there are less than $6(g+m)$ edges in $T^{1}$ by Lemma C.2.
@@ C.2 Analysis of DeleteVertices
Proof C.7.
Let $m_{1}$ and $m_{2}$ count respectively the edges and the faces of $Y$. Euler’s formula gives $6m-6m_{1}+6m_{2}=12-12g-6b$. Double counting gives $3m_{2}\leq 2m_{1}-b$ and $2m_{1}=\sum_{v}\deg v$, where the sum is over the vertices, and where $\deg v$ denotes the degree of a vertex $v$. Then $\sum_{v}6-\deg v=6m-2m_{1}\geq 6m-6m_{1}+6m_{2}+2b\geq 12-12g-4b>-m/2$. Let $a$ and $b$ count the number of vertices whose degree is respectively smaller than or equal to six, and greater than six. Then $b<5a+m/2$. Assuming $a<m/12$, we get $b<11m/12$, and so $a+b<m$. This is a contradiction. This proves the lemma.
Proof C.8 (Proof of Lemma C.5).
Let $b$ be the number of boundary components of $\mathcal{S}(T)$. We have $m>24(g+b)$. Indeed we assumed $m>24(g+\bar{m})$, and we have $\bar{m}\geq b$ as every boundary component of $\mathcal{S}(T)$ contains a strong vertex of $T^{1}$. So by Lemma C.6 at least $m/12$ vertices of $T^{1}$ have degree smaller than or equal to six. Moreover less than $m/24$ vertices of $T^{1}$ are strong by assumption. So more than $m/24$ vertices of $T^{1}$ are weak and have degree smaller than or equal to six. Any maximal independent set of such vertices contains more than $m/(24\times 7)=m/168$ vertices, so DeleteVertices deletes more than $m/168$ vertices.
@@ C.3 Analysis of SimplifyTubes
Now for every $T_{0}\in Z^{\prime}$ let $b(T_{0})$ be the number of boundary components of $\mathcal{S}(T_{0})$. The claim implies $b(T_{0})\leq 2-\chi(T_{0})\leq 2+\lambda(T_{0})\leq 3\lambda(T_{0})$. So $\sum_{T_{0}\in Z^{\prime}}b(T_{0})\leq 3\sum_{T_{0}\in Z^{\prime}}\lambda(T_{0})\leq 3\sum_{T_{0}\in Z}\lambda(T_{0})\leq 9(g+b)$. Therefore at most $9(g+b)$ loops in $I$ are incident to the surface of some $T_{0}\in Z^{\prime}$. If every other loop in $I$ is incident to the surfaces of two distinct $T_{0},T_{1}\in Z$ then we are done. Otherwise there is a loop $e\in I$ incident to the surface of only one $T_{0}\in Z$. Because $T_{0}$ is a tube, $\mathcal{S}(T)$ is a homeomorphic to a torus, and $e$ is the only loop in $I$, so we are done. This proves the lemma.
Proof C.12 (Proof of Lemma C.9).
@@
In the particular case where the surface $\mathcal{S}(T)$ is homeomorphic to a torus, and where $T^{1}$ contains exactly one vertex incident to loop a edge, the application of SimplifyTubes does nothing and $T=T^{\prime}$. In this case the lemma is proved. In all other cases if a vertex $v$ of $T^{\prime 1}$ lies in the interior of $\mathcal{S}(T^{\prime})$ and is incident to a loop edge in $T^{\prime 1}$, then $v$ is the base vertex of some loop in $J$. Indeed $v$ would otherwise have been deleted by SimplifyTubes when replacing a tube by a biface. There are at most $9(g+b)$ such vertices by our claim. This proves the lemma.
@@ C.4 Analysis of Gardening
Proof C.14.
The genus of $\mathcal{S}(R_{A})$ is smaller than or equal to $g$. It is the number of boundary components of $\mathcal{S}(R_{A})$ that we must handle. Each boundary component of $\mathcal{S}(R_{A})$ is either a boundary component of $\mathcal{S}(R)$, and there are $b$ of them, or it is a separating loop. We bound the the number of separating loops adjacent to $\mathcal{S}(R_{A})$, so let $I$ contain those loops. Each $e\in I$ is incident to two connected components of $\mathcal{S}(R)\setminus I$: one of them is in $\mathcal{S}(R_{A})$, the other is not. The component in $\mathcal{S}(R_{A})$ is not the surface of a tube because Gardening was just applied. So $I$ contains at most $9(g+b)$ loops by Lemma C.10. We proved that $\mathcal{S}(R_{A})$ has at most $10(g+b)$ boundary components.
@@
Now we prove the lemma by considering a single iteration of the main loop. Assuming that $R_{A}^{1}$ has more than $3000(g+b+m)$ vertices at the beginning of the iteration, we shall prove that in the end of the iteration the number of vertices of $R_{A}^{1}$ has decreased. To do so first observe that the iteration starts with InsertVertices, and this is the only moment where vertices are inserted. At this point the number of vertices of $R_{A}^{1}$ is multiplied by less than $8$ by Lemma C.1. And by our claim, as long as the number of vertices exceeds $3000(g+b+m)$ it is divided by more than $168/167$ by each iteration of the interior loop. There are 350 iterations of the interior loop, and $8<(168/167)^{350}$. This proves the lemma.
@@
See 5.9
Proof D.1 (Proof of Lemma 5.9).
Let $t>1$. Assume that there is a loop $\gamma$, based at a point $x$, that encloses $f$ by a factor of $t$. Then $\gamma$ encloses $e$ by a factor of $t$ because $\langle x\rangle_{f}\leq\langle x\rangle_{e}$.
@@
We have $\ell(\gamma)<\langle y\rangle_{f}$ so $\gamma$ is in general position with $f$. Direct $\gamma$ so that $\gamma$ crosses $f$ from right to left at $y$, and write $\gamma$ as the concatenation of two paths $\gamma_{0}$ and $\gamma_{1}$ respectively before and after its crossing at $y$. There is a lift $\widetilde{\gamma}_{1}$ of $\gamma_{1}$ that leaves $\widetilde{y}$ on the left of $\widetilde{f}$. And $\widetilde{\gamma}_{1}$ is otherwise disjoint from $\widetilde{L}$, because the interior of $\widetilde{S}$ is flat. Thus the endpoint $\widetilde{x}$ of $\widetilde{\gamma}_{1}$ lies on the left of $\widetilde{L}$. There is a lift $\widetilde{\gamma}_{0}$ of $\gamma_{0}$ that starts at $\widetilde{x}$. And $\widetilde{\gamma}_{0}$ is otherwise disjoint from $\widetilde{\gamma}_{1}$ because $\gamma$ meets $x$ on both sides of $e$, and because the interior of $\widetilde{S}$ is flat. By the previous paragraph, the endpoint of $\widetilde{\gamma}_{0}$ lies on the right side of $\widetilde{L}$, so $\widetilde{\gamma}_{0}$ intersects $\widetilde{L}$. Cut $\widetilde{\gamma}_{0}$ at its first intersection point $\widetilde{z}$ with $\widetilde{L}$. Let $\widetilde{I}$ be the sub-segment of $\widetilde{L}$ between $\widetilde{y}$ and $\widetilde{z}$. The concatenation of the prefix of $\widetilde{\gamma}_{0}$ ending at $\widetilde{z}$, of $\widetilde{I}$, and of $\widetilde{\gamma}_{1}$ is a simple closed curve $\widetilde{C}$. At $\widetilde{x}$, there is a portion of $\widetilde{e}$ that enters the bounded side of $\widetilde{C}$, because $\gamma$ meets $x$ on both sides of $e$. This portion of $\widetilde{e}$ can be extended into a geodesic $\widetilde{p}$ that meets $\widetilde{C}$ at some point $\widetilde{v}$, because the interior of $\widetilde{S}$ is flat. Then $\widetilde{v}$ belongs to the relative interior of $\widetilde{I}$. We claim that $\widetilde{v}$ belongs to the relative interiors of both $\widetilde{e}$ and $\widetilde{f}$, which is a contradiction because the relative interiors of $e$ and $f$ are disjoint. To prove the claim, first observe that the distance between $\widetilde{y}$ and $\widetilde{z}$ in $\widetilde{S}$ is at most $\ell(\gamma)$, and this distance is equal to the length of $\widetilde{I}$, because the interior of $\widetilde{S}$ is flat. So the sub-segment of $\widetilde{I}$ between $\widetilde{y}$ and $\widetilde{v}$ is no longer than $\ell(\gamma)<\langle y\rangle_{f}$, and is thus included in the relative interior of $\widetilde{f}$. Also, the distance between $\widetilde{v}$ and $\widetilde{x}$ is smaller than or equal to $2\ell(\gamma)\leq 2\langle x\rangle_{e}/t<\langle x\rangle_{e}$, so $\widetilde{p}$ is included in the relative interior of $\widetilde{e}$.
Figure 5: The setting of Lemma D.4.
@@
The proof of Proposition 5.12 also relies on the following construction. See Figure 5. In the Euclidean plane $\mathbb{R}^{2}$ let $Q$ be a polygon with more than three vertices. Let $e$ be a shortcut of $Q$. Let $f$ and $f^{\prime}$ be sides of $Q$ separated by $e$ along the boundary of $Q$. Let $x$ be a point in the relative interior of $e$. Let $y$ and $y^{\prime}$ be points that lie on respectively $f$ and $f^{\prime}$ (possibly vertices of $Q$), and do not lie on $e$. Consider the segments $p$ and $p^{\prime}$ between $x$ and respectively $y$ and $y^{\prime}$. Assume that the relative interiors $p$ and $p^{\prime}$ are included in the interior of $Q$. Then:
@@
Applying the algorithm to $p^{\prime}$ and $f^{\prime}$ on the other side of $e$, either the algorithm does not break in which case $\ell(f^{\prime})\geq(1-4/t)\ell(e)$, $\langle y^{\prime}\rangle_{f^{\prime}}\geq(1-4/t)\langle x\rangle_{e}$, and we are done. Or the algorithm breaks and we get similarly a triangle $\Delta^{\prime}$ and a vertex $w^{\prime}$ of $P$. The inner angles of $\Delta^{\prime}$ at $v_{0}$ and $v_{1}$ are also both strictly smaller than $\pi/4$, so the relative interior of the segment between $w$ and $w^{\prime}$ is included in the interior of the quadrilateral formed by $\Delta$ and $\Delta^{\prime}$, and is strictly shorter than $e$. This segment is a vertex-to-vertex arc of $Q$ shorter than $e$, a contradiction.
Proof D.6 (Proof of Proposition 5.12).
@@ D.3 Interior of a thick biface
Proof D.9.
Among the two interior edges of $B^{1}$ let $e$ be a shortest one, and let $g\neq e$ be the other one. Then $e$, $g$, and $f$ are the sides of $F$. The angle at the corner of $F$ between $f$ and $g$ is smaller than $\pi/2$ because $\ell(e)\leq\ell(g)$. Now consider the corner $c$ between $f$ and $e$. Cut $\mathcal{S}(B)$ open along $e$ and consider the resulting quadrilateral $Q$ in the plane. The edge $f$ of $B^{1}$ corresponds to a side $\widehat{f}$ of $Q$, the edge $e$ corresponds to two opposite sides $\widehat{e}$ and $\widehat{e}^{\prime}$, and the edge $g$ corresponds to a vertex-to-vertex arc $\widehat{g}$ of $Q$. Also the other boundary edge $f^{\prime}\neq f$ of $B^{1}$ corresponds to the side $\widehat{f}^{\prime}$ of $Q$ opposite to $\widehat{f}$. And the corner $c$ corresponds to the corner $\widehat{c}$ of $Q$ between $\widehat{e}$ and $\widehat{f}$. Let $\widehat{d}$ be the corner of $Q$ opposite to $\widehat{c}$, between $\widehat{e}^{\prime}$ and $\widehat{f}^{\prime}$. Assume by contradiction that the angle at $\widehat{c}$ is greater than $\pi/2$. We have $\ell(\widehat{e})=\ell(\widehat{e}^{\prime})$ and $\ell(\widehat{f})\geq\ell(\widehat{f}^{\prime})$ so the angle at $\widehat{d}$ is greater than or equal to the angle at $\widehat{c}$, and in particular is also greater than $\pi/2$. The two other angles of $Q$ are smaller than $\pi$, so $Q$ is convex and admits a diagonal $p\neq\widehat{g}$. Consider the unique circle $C$ that admits $\widehat{g}$ as a diameter. Then the two endpoints of $p$ lie in the interior of $C$. So $p$ is shorter than $\widehat{g}$. This contradicts the assumption that $B$ is good.
Proof D.10 (Proof of Proposition D.7).
@@
If $e=g$ we are done by our claim, so assume that $e$ is a longest interior edge of $B^{1}$. Deleting $e$ merges the two faces of $B^{1}$ into a single face $F^{\prime}$ of which $e$ is a shortcut, because $B$ is good. So Proposition 5.12 applies because $c_{S}(e)>6$: there is a boundary edge $f$ of $F^{\prime}$ such that $c_{S}(f)\geq c_{S}(e)-4$ and $\ell(f)\geq(1-4/c_{S}(e))\ell(e)$. If $f$ is a boundary edge of $B^{1}$ we are done. Otherwise $f=g$ so $\ell(g^{\prime})\geq\ell(f)\geq(1-4/c_{S}(e))\ell(e)$ and $c_{S}(g^{\prime})\geq c_{S}(f)-1\geq c_{S}(e)-5$ by our claim because $c_{S}(e)>6$. This proves the proposition.
@@
Proof D.12.
Assume by contradiction that there is a corner $c$ between $e$ and a boundary edge $f$ of $B^{1}$ whose angle is smaller than or equal to $\pi/4$. Cut $\mathcal{S}(B)$ open along $e$ and embed the resulting quadrilateral $Q$ in the plane, isometrically. The edge $e$ corresponds to two opposite sides $\widehat{e}$ and $\widehat{e}^{\prime}$ of $Q$. The edge $f$ corresponds to one of the other two sides of $Q$, that we call $\widehat{f}$. The vertex $v$ of the corner $c$ corresponds to the two end-vertices of $\widehat{f}$: let $\widehat{v}$ be the one incident to $\widehat{e}$, and let $\widehat{v}^{\prime}$ be the one incident to $\widehat{e}^{\prime}$. Without loss of generality the corner $c$ corresponds to the corner of $Q$ at $\widehat{v}$, whose angle is thus smaller than or equal to $\pi/4$. Consider the orthogonal projection $x$ of $\widehat{v}^{\prime}$ on the line containing $\widehat{e}$. Then $x$ belongs to $\widehat{e}$ because $\widehat{e}$ is longer than $\widehat{f}$, as $B$ is thin. The segment $p$ between $x$ and $\widehat{v}^{\prime}$ is shorter than the portion of $\widehat{e}$ between $x$ and $\widehat{v}$. Also $p$ is included in $Q$ because $\widehat{e}$ and $\widehat{e}^{\prime}$ are longer than $\widehat{f}$. Thus $p$ projects to a path that shortcuts $e$, contradicting the fact that $B$ is a good biface.
@@
Proof D.14.
Without loss of generality one of the two endpoints of $p$ (at least) is a vertex $v$ of $B^{1}$. Consider the other endpoint $x$ of $p$, and the vertex $w\neq v$ of $B^{1}$. There is a path $q$ from $x$ to $w$ in the boundary of $\mathcal{S}(B)$. Without loss of generality $\ell(q)\leq\ell(e)/2$ because $B$ is thin. Also $e$ is a shortest path because $B$ is good. So $\ell(p)+\ell(q)\geq\ell(e)$. We proved $\ell(p)\geq\ell(e)/2$.
Proof D.15 (Proof of Proposition 5.5).
Among the two interior edges of $B^{1}$ let $e$ be a shortest one. Let $f$ be any one of the two boundary edges of $B^{1}$. We have $\ell(e)\geq\ell(f)$ because $B$ is thin. Assume by contradiction that there is in $S$ a loop $\gamma$ that encloses $f$ by a factor of $t>2$. Let $x$ be the basepoint of $\gamma$. There is a portion of $\gamma$ that leaves $x$ and enters the interior of $\mathcal{S}(B)$. This portion of $\gamma$ cannot leave $\mathcal{S}(B)$ via the other boundary edge of $\mathcal{S}(B)$, for otherwise $\ell(\gamma)\geq\ell(e)/2$ by Lemma D.13, so $\ell(\gamma)>\langle x\rangle_{f}/t$, a contradiction. Then $\gamma$ intersects $e$. And $f$ and $e$ have a corner whose angle is smaller than $\pi/4$ because $\ell(\gamma)<\langle x\rangle_{f}/2$. This contradicts Lemma D.11.
@@
To prove that $\widetilde{a}_{1}$ lies strictly to the right of $\widetilde{L}$, consider the lift $\widetilde{y}$ of $y$ in $\widetilde{e}$, and the lift $\widetilde{p}_{1}$ of $p$ that ends at $\widetilde{y}$. Then the startpoint of $\widetilde{p}_{1}$ is $\widetilde{a}_{1}$, and it lies strictly to the right of $\widetilde{L}$ because $\widetilde{p}_{1}$ meets $\widetilde{y}$ on the right of $\widetilde{L}$, and because $\widetilde{p}_{1}$ is otherwise disjoint from $\widetilde{L}$. This proves the claim, and the lemma.

### v2

@@ C.1 Analysis of InsertVertices
Proof.
Let $m_{1}$ and $m_{2}$ count respectively the edges and the faces of $T^{1}$, and let $b$ count the boundary components of $\mathcal{S}(T)$. Double counting gives $3m_{2}\leq 2m_{1}$. Euler’s formula gives $m_{1}-m_{2}=m+2g+b-2$. And we have $b\leq m$. Therefore $m_{1}\leq 3m_{1}-3m_{2}<6(m+g)$. ∎
Proof of Lemma C.1.
There are no more vertices inserted than there are edges in $T^{1}$, and there are less than $6(g+m)$ edges in $T^{1}$ by Lemma C.2. ∎
@@ C.2 Analysis of DeleteVertices
Proof.
Let $m_{1}$ and $m_{2}$ count respectively the edges and the faces of $Y$. Euler’s formula gives $6m-6m_{1}+6m_{2}=12-12g-6b$. Double counting gives $3m_{2}\leq 2m_{1}-b$ and $2m_{1}=\sum_{v}\deg v$, where the sum is over the vertices, and where $\deg v$ denotes the degree of a vertex $v$. Then $\sum_{v}6-\deg v=6m-2m_{1}\geq 6m-6m_{1}+6m_{2}+2b\geq 12-12g-4b>-m/2$. Let $a$ and $b$ count the number of vertices whose degree is respectively smaller than or equal to six, and greater than six. Then $b<5a+m/2$. Assuming $a<m/12$, we get $b<11m/12$, and so $a+b<m$. This is a contradiction. This proves the lemma. ∎
Proof of Lemma C.3.
Let $b$ be the number of boundary components of $\mathcal{S}(T)$. We have $m>24(g+b)$. Indeed we assumed $m>24(g+\bar{m})$, and we have $\bar{m}\geq b$ as every boundary component of $\mathcal{S}(T)$ contains a strong vertex of $T^{1}$. So by Lemma C.4 at least $m/12$ vertices of $T^{1}$ have degree smaller than or equal to six. Moreover less than $m/24$ vertices of $T^{1}$ are strong by assumption. So more than $m/24$ vertices of $T^{1}$ are weak and have degree smaller than or equal to six. Any maximal independent set of such vertices contains more than $m/(24\times 7)=m/168$ vertices, so DeleteVertices deletes more than $m/168$ vertices. ∎
@@ C.3 Analysis of SimplifyTubes
Now for every $T_{0}\in Z^{\prime}$ let $b(T_{0})$ be the number of boundary components of $\mathcal{S}(T_{0})$. The claim implies $b(T_{0})\leq 2-\chi(T_{0})\leq 2+\lambda(T_{0})\leq 3\lambda(T_{0})$. So $\sum_{T_{0}\in Z^{\prime}}b(T_{0})\leq 3\sum_{T_{0}\in Z^{\prime}}\lambda(T_{0})\leq 3\sum_{T_{0}\in Z}\lambda(T_{0})\leq 9(g+b)$. Therefore at most $9(g+b)$ loops in $I$ are incident to the surface of some $T_{0}\in Z^{\prime}$. If every other loop in $I$ is incident to the surfaces of two distinct $T_{0},T_{1}\in Z$ then we are done. Otherwise there is a loop $e\in I$ incident to the surface of only one $T_{0}\in Z$. Because $T_{0}$ is a tube, $\mathcal{S}(T)$ is a homeomorphic to a torus, and $e$ is the only loop in $I$, so we are done. This proves the lemma. ∎
Proof of Lemma C.5.
@@
In the particular case where the surface $\mathcal{S}(T)$ is homeomorphic to a torus, and where $T^{1}$ contains exactly one vertex incident to loop a edge, the application of SimplifyTubes does nothing and $T=T^{\prime}$. In this case the lemma is proved. In all other cases if a vertex $v$ of $T^{\prime 1}$ lies in the interior of $\mathcal{S}(T^{\prime})$ and is incident to a loop edge in $T^{\prime 1}$, then $v$ is the base vertex of some loop in $J$. Indeed $v$ would otherwise have been deleted by SimplifyTubes when replacing a tube by a biface. There are at most $9(g+b)$ such vertices by our claim. This proves the lemma. ∎
@@ C.4 Analysis of Gardening
Proof.
The genus of $\mathcal{S}(R_{A})$ is smaller than or equal to $g$. It is the number of boundary components of $\mathcal{S}(R_{A})$ that we must handle. Each boundary component of $\mathcal{S}(R_{A})$ is either a boundary component of $\mathcal{S}(R)$, and there are $b$ of them, or it is a separating loop. We bound the the number of separating loops adjacent to $\mathcal{S}(R_{A})$, so let $I$ contain those loops. Each $e\in I$ is incident to two connected components of $\mathcal{S}(R)\setminus I$: one of them is in $\mathcal{S}(R_{A})$, the other is not. The component in $\mathcal{S}(R_{A})$ is not the surface of a tube because Gardening was just applied. So $I$ contains at most $9(g+b)$ loops by Lemma C.6. We proved that $\mathcal{S}(R_{A})$ has at most $10(g+b)$ boundary components. ∎
@@
Now we prove the lemma by considering a single iteration of the main loop. Assuming that $R_{A}^{1}$ has more than $3000(g+b+m)$ vertices at the beginning of the iteration, we shall prove that in the end of the iteration the number of vertices of $R_{A}^{1}$ has decreased. To do so first observe that the iteration starts with InsertVertices, and this is the only moment where vertices are inserted. At this point the number of vertices of $R_{A}^{1}$ is multiplied by less than $8$ by Lemma C.1. And by our claim, as long as the number of vertices exceeds $3000(g+b+m)$ it is divided by more than $168/167$ by each iteration of the interior loop. There are 350 iterations of the interior loop, and $8<(168/167)^{350}$. This proves the lemma. ∎
@@
See 5.7
Proof of Lemma 5.7.
Let $t>1$. Assume that there is a loop $\gamma$, based at a point $x$, that encloses $f$ by a factor of $t$. Then $\gamma$ encloses $e$ by a factor of $t$ because $\langle x\rangle_{f}\leq\langle x\rangle_{e}$. ∎
@@
We have $\ell(\gamma)<\langle y\rangle_{f}$ so $\gamma$ is in general position with $f$. Direct $\gamma$ so that $\gamma$ crosses $f$ from right to left at $y$, and write $\gamma$ as the concatenation of two paths $\gamma_{0}$ and $\gamma_{1}$ respectively before and after its crossing at $y$. There is a lift $\widetilde{\gamma}_{1}$ of $\gamma_{1}$ that leaves $\widetilde{y}$ on the left of $\widetilde{f}$. And $\widetilde{\gamma}_{1}$ is otherwise disjoint from $\widetilde{L}$, because the interior of $\widetilde{S}$ is flat. Thus the endpoint $\widetilde{x}$ of $\widetilde{\gamma}_{1}$ lies on the left of $\widetilde{L}$. There is a lift $\widetilde{\gamma}_{0}$ of $\gamma_{0}$ that starts at $\widetilde{x}$. And $\widetilde{\gamma}_{0}$ is otherwise disjoint from $\widetilde{\gamma}_{1}$ because $\gamma$ meets $x$ on both sides of $e$, and because the interior of $\widetilde{S}$ is flat. By the previous paragraph, the endpoint of $\widetilde{\gamma}_{0}$ lies on the right side of $\widetilde{L}$, so $\widetilde{\gamma}_{0}$ intersects $\widetilde{L}$. Cut $\widetilde{\gamma}_{0}$ at its first intersection point $\widetilde{z}$ with $\widetilde{L}$. Let $\widetilde{I}$ be the sub-segment of $\widetilde{L}$ between $\widetilde{y}$ and $\widetilde{z}$. The concatenation of the prefix of $\widetilde{\gamma}_{0}$ ending at $\widetilde{z}$, of $\widetilde{I}$, and of $\widetilde{\gamma}_{1}$ is a simple closed curve $\widetilde{C}$. At $\widetilde{x}$, there is a portion of $\widetilde{e}$ that enters the bounded side of $\widetilde{C}$, because $\gamma$ meets $x$ on both sides of $e$. This portion of $\widetilde{e}$ can be extended into a geodesic $\widetilde{p}$ that meets $\widetilde{C}$ at some point $\widetilde{v}$, because the interior of $\widetilde{S}$ is flat. Then $\widetilde{v}$ belongs to the relative interior of $\widetilde{I}$. We claim that $\widetilde{v}$ belongs to the relative interiors of both $\widetilde{e}$ and $\widetilde{f}$, which is a contradiction because the relative interiors of $e$ and $f$ are disjoint. To prove the claim, first observe that the distance between $\widetilde{y}$ and $\widetilde{z}$ in $\widetilde{S}$ is at most $\ell(\gamma)$, and this distance is equal to the length of $\widetilde{I}$, because the interior of $\widetilde{S}$ is flat. So the sub-segment of $\widetilde{I}$ between $\widetilde{y}$ and $\widetilde{v}$ is no longer than $\ell(\gamma)<\langle y\rangle_{f}$, and is thus included in the relative interior of $\widetilde{f}$. Also, the distance between $\widetilde{v}$ and $\widetilde{x}$ is smaller than or equal to $2\ell(\gamma)\leq 2\langle x\rangle_{e}/t<\langle x\rangle_{e}$, so $\widetilde{p}$ is included in the relative interior of $\widetilde{e}$. ∎
Figure D.1: The setting of Lemma D.2.
@@
The proof of Proposition 5.9 also relies on the following construction. See Figure D.1. In the Euclidean plane $\mathbb{R}^{2}$ let $Q$ be a polygon with more than three vertices. Let $e$ be a shortcut of $Q$. Let $f$ and $f^{\prime}$ be sides of $Q$ separated by $e$ along the boundary of $Q$. Let $x$ be a point in the relative interior of $e$. Let $y$ and $y^{\prime}$ be points that lie on respectively $f$ and $f^{\prime}$ (possibly vertices of $Q$), and do not lie on $e$. Consider the segments $p$ and $p^{\prime}$ between $x$ and respectively $y$ and $y^{\prime}$. Assume that the relative interiors $p$ and $p^{\prime}$ are included in the interior of $Q$. Then:
@@
Applying the algorithm to $p^{\prime}$ and $f^{\prime}$ on the other side of $e$, either the algorithm does not break in which case $\ell(f^{\prime})\geq(1-4/t)\ell(e)$, $\langle y^{\prime}\rangle_{f^{\prime}}\geq(1-4/t)\langle x\rangle_{e}$, and we are done. Or the algorithm breaks and we get similarly a triangle $\Delta^{\prime}$ and a vertex $w^{\prime}$ of $P$. The inner angles of $\Delta^{\prime}$ at $v_{0}$ and $v_{1}$ are also both strictly smaller than $\pi/4$, so the relative interior of the segment between $w$ and $w^{\prime}$ is included in the interior of the quadrilateral formed by $\Delta$ and $\Delta^{\prime}$, and is strictly shorter than $e$. This segment is a vertex-to-vertex arc of $Q$ shorter than $e$, a contradiction. ∎
Proof of Proposition 5.9.
@@ D.3 Interior of a thick biface
Proof.
Among the two interior edges of $B^{1}$ let $e$ be a shortest one, and let $g\neq e$ be the other one. Then $e$, $g$, and $f$ are the sides of $F$. The angle at the corner of $F$ between $f$ and $g$ is smaller than $\pi/2$ because $\ell(e)\leq\ell(g)$. Now consider the corner $c$ between $f$ and $e$. Cut $\mathcal{S}(B)$ open along $e$ and consider the resulting quadrilateral $Q$ in the plane. The edge $f$ of $B^{1}$ corresponds to a side $\widehat{f}$ of $Q$, the edge $e$ corresponds to two opposite sides $\widehat{e}$ and $\widehat{e}^{\prime}$, and the edge $g$ corresponds to a vertex-to-vertex arc $\widehat{g}$ of $Q$. Also the other boundary edge $f^{\prime}\neq f$ of $B^{1}$ corresponds to the side $\widehat{f}^{\prime}$ of $Q$ opposite to $\widehat{f}$. And the corner $c$ corresponds to the corner $\widehat{c}$ of $Q$ between $\widehat{e}$ and $\widehat{f}$. Let $\widehat{d}$ be the corner of $Q$ opposite to $\widehat{c}$, between $\widehat{e}^{\prime}$ and $\widehat{f}^{\prime}$. Assume by contradiction that the angle at $\widehat{c}$ is greater than $\pi/2$. We have $\ell(\widehat{e})=\ell(\widehat{e}^{\prime})$ and $\ell(\widehat{f})\geq\ell(\widehat{f}^{\prime})$ so the angle at $\widehat{d}$ is greater than or equal to the angle at $\widehat{c}$, and in particular is also greater than $\pi/2$. The two other angles of $Q$ are smaller than $\pi$, so $Q$ is convex and admits a diagonal $p\neq\widehat{g}$. Consider the unique circle $C$ that admits $\widehat{g}$ as a diameter. Then the two endpoints of $p$ lie in the interior of $C$. So $p$ is shorter than $\widehat{g}$. This contradicts the assumption that $B$ is good. ∎
Proof of Proposition D.3.
@@
If $e=g$ we are done by our claim, so assume that $e$ is a longest interior edge of $B^{1}$. Deleting $e$ merges the two faces of $B^{1}$ into a single face $F^{\prime}$ of which $e$ is a shortcut, because $B$ is good. So Proposition 5.9 applies because $c_{S}(e)>6$: there is a boundary edge $f$ of $F^{\prime}$ such that $c_{S}(f)\geq c_{S}(e)-4$ and $\ell(f)\geq(1-4/c_{S}(e))\ell(e)$. If $f$ is a boundary edge of $B^{1}$ we are done. Otherwise $f=g$ so $\ell(g^{\prime})\geq\ell(f)\geq(1-4/c_{S}(e))\ell(e)$ and $c_{S}(g^{\prime})\geq c_{S}(f)-1\geq c_{S}(e)-5$ by our claim because $c_{S}(e)>6$. This proves the proposition. ∎
@@
Proof.
Assume by contradiction that there is a corner $c$ between $e$ and a boundary edge $f$ of $B^{1}$ whose angle is smaller than or equal to $\pi/4$. Cut $\mathcal{S}(B)$ open along $e$ and embed the resulting quadrilateral $Q$ in the plane, isometrically. The edge $e$ corresponds to two opposite sides $\widehat{e}$ and $\widehat{e}^{\prime}$ of $Q$. The edge $f$ corresponds to one of the other two sides of $Q$, that we call $\widehat{f}$. The vertex $v$ of the corner $c$ corresponds to the two end-vertices of $\widehat{f}$: let $\widehat{v}$ be the one incident to $\widehat{e}$, and let $\widehat{v}^{\prime}$ be the one incident to $\widehat{e}^{\prime}$. Without loss of generality the corner $c$ corresponds to the corner of $Q$ at $\widehat{v}$, whose angle is thus smaller than or equal to $\pi/4$. Consider the orthogonal projection $x$ of $\widehat{v}^{\prime}$ on the line containing $\widehat{e}$. Then $x$ belongs to $\widehat{e}$ because $\widehat{e}$ is longer than $\widehat{f}$, as $B$ is thin. The segment $p$ between $x$ and $\widehat{v}^{\prime}$ is shorter than the portion of $\widehat{e}$ between $x$ and $\widehat{v}$. Also $p$ is included in $Q$ because $\widehat{e}$ and $\widehat{e}^{\prime}$ are longer than $\widehat{f}$. Thus $p$ projects to a path that shortcuts $e$, contradicting the fact that $B$ is a good biface. ∎
@@
Proof.
Without loss of generality one of the two endpoints of $p$ (at least) is a vertex $v$ of $B^{1}$. Consider the other endpoint $x$ of $p$, and the vertex $w\neq v$ of $B^{1}$. There is a path $q$ from $x$ to $w$ in the boundary of $\mathcal{S}(B)$. Without loss of generality $\ell(q)\leq\ell(e)/2$ because $B$ is thin. Also $e$ is a shortest path because $B$ is good. So $\ell(p)+\ell(q)\geq\ell(e)$. We proved $\ell(p)\geq\ell(e)/2$. ∎
Proof of Proposition 5.4.
Among the two interior edges of $B^{1}$ let $e$ be a shortest one. Let $f$ be any one of the two boundary edges of $B^{1}$. We have $\ell(e)\geq\ell(f)$ because $B$ is thin. Assume by contradiction that there is in $S$ a loop $\gamma$ that encloses $f$ by a factor of $t>2$. Let $x$ be the basepoint of $\gamma$. There is a portion of $\gamma$ that leaves $x$ and enters the interior of $\mathcal{S}(B)$. This portion of $\gamma$ cannot leave $\mathcal{S}(B)$ via the other boundary edge of $\mathcal{S}(B)$, for otherwise $\ell(\gamma)\geq\ell(e)/2$ by Lemma D.6, so $\ell(\gamma)>\langle x\rangle_{f}/t$, a contradiction. Then $\gamma$ intersects $e$. And $f$ and $e$ have a corner whose angle is smaller than $\pi/4$ because $\ell(\gamma)<\langle x\rangle_{f}/2$. This contradicts Lemma D.5. ∎
@@
To prove that $\widetilde{a}_{1}$ lies strictly to the right of $\widetilde{L}$, consider the lift $\widetilde{y}$ of $y$ in $\widetilde{e}$, and the lift $\widetilde{p}_{1}$ of $p$ that ends at $\widetilde{y}$. Then the startpoint of $\widetilde{p}_{1}$ is $\widetilde{a}_{1}$, and it lies strictly to the right of $\widetilde{L}$ because $\widetilde{p}_{1}$ meets $\widetilde{y}$ on the right of $\widetilde{L}$, and because $\widetilde{p}_{1}$ is otherwise disjoint from $\widetilde{L}$. This proves the claim, and the lemma. ∎